OpenStudy (anonymous):

how do I simplify: x^2+5x+4=x^2+2x-15?

OpenStudy (anonymous):

3x + 19 = 0 if u need to solve it x = -19/3

OpenStudy (anonymous):

I do not understand how 3x was gotten though..

OpenStudy (turingtest):

5x-2x

OpenStudy (anonymous):

by subtractin 2x from both sides

OpenStudy (anonymous):

The x^2 on both sides of the equation subtract out. 5x + 4 = 2x - 15 3x = -19 x = -19/3

OpenStudy (anonymous):

thank you very much, lol I was just about to ask how to get rid of the x^2

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

ty for the explanation